Using the 5 step process; solve the following problem:
The sum of the page numbers on the facing pages of a book is 285. Find the larger page number.

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

let one page number be x
the other page will be x+1
..
x+x+1=285
2x+1=285
add-1 to both sides
2x+1-1 =285-1
2x=284
divide by 2
2x/2 = 284/2
x=142
the other page number is 143

Answer by Alan3354(69443)

285/2 = 142+
--> 142 & 143